OpenStudy (anonymous):

8m+7=5m+16 show work

OpenStudy (anonymous):

3m = 9 m = 3

OpenStudy (anonymous):

Take 5m from both sides of the equation, then take 7 from both sides of the equation. What does that give you?

jhonyy9 (jhonyy9):

this is right ?

jhonyy9 (jhonyy9):

you need to do what have wrote there

OpenStudy (katielong):

8m+7=5m+16 -7 8m=5m+16-7 simplify 8m=5m+9 -5m 8m-5m=9 simplify 3m=9 soo.... m=3 simples